# Theory of two-component superfluidity of microcavity polaritons

### v1

@@ V Superfluidity
where $T_{c}$ is the critical temperature for superfluidity, $\hbar$ is the reduced Planck’s constant, $\rho$ is the total 2D density of the system, accounting for contributions from both polariton branches, $c_{s}$ is the sound velocity defined for the two-component condensate in Eq.  (39) and for the one-component condensate in Eq. (40). Additionally, $\zeta(3)$ is the Riemann zeta function evaluated for a value of $3$, $s$ is the spin degeneracy factor and $k_{B}$ is the Boltzmann constant. For the two component condensate of upper and lower polaritons, this is explicitly written in terms of the Hopfield coefficients and effective masses of the upper and lower polartions as follows
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho}{3\zeta(3)sk_{B}^{3}}\right]^{1/3}\left({U_{q}n_{0}\left(\frac{|X_{P}|^{2}}{m_{\text{LP}}}+\frac{|C_{P}|^{2}}{m_{\text{UP}}}\right)}\right)^{2/3}.$ (45)
For the two component superfluid, we emphasise that $\rho=m_{\text{UP}}n_{0}|C_{P}|^{2}+m_{\text{LP}}n_{0}|X_{P}|^{2}$ is the total 2D density of the system. For comparison we will also be studying how this fares compared to the critical temperature for a one-component superfluid of only lower polaritons
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho}{3\,\zeta(3)\,s\,k_{B}^{3}}\right]^{{1}/{3}}\left(U_{q}\,n_{0}\,\frac{|X_{P}|^{2}}{m_{\mathrm{LP}}}\right)^{2/3}.$ (46)
For this case, it is worth noting that while $n_{0}$ is still the total concentration of polaritons, only the lower polaritons can form a condensate and as such $\rho=m_{\text{LP}}n_{0}|X_{P}|^{2}$ is the correct density to use. We plot the dependance of the critical temperature from Eq. (45) as a function of the detuning in Fig. 7 .
Figure 7: The critical temperature $T_{c}$ as a function of detuning in GaAs for different Rabi splittings. The plots are symmetric with respect to detuning away from resonance, meaning the values for positive and negative detuning (e.g., $\Delta=10\,\text{meV}$ and $\Delta=-10\,\text{meV}$) are identical. The parameters used are a total polariton concentration of $n_{0}=10^{10}\,\text{cm}^{-2}$ and a bare photon energy of  1.6 eV. The critical temperature was calculated using the definition from Eq. (45).
The plot highlights the symmetric nature of $T_{c}$ with respect to $\Delta_{0}=0$, meaning that the critical temperature profiles for positive and negative detuning are identical. The maximum $T_{c}$ occurs at resonance ($\Delta_{0}=0$), where the polaritonic system is most balanced between its excitonic and photonic components. The critical temperature is fairly stable around zero detuning, and exhibits a plateau around that region.
Detuning away from resonance generally leads to a lower critical temperatures. This trend reflects the increased dominance of either the excitonic or photonic fraction in the polariton composition, which causes a lower $T_{c}$. However, the Rabi splitting allows us to change the critical temperature. A larger Rabi splitting leads to larger critical temperature for all materials, across all non-zero detuning levels.
Figure 8: (COLOUR ONLINE) Critical temperature ($T_{c}$) as a function of detuning ($\Delta_{0}$) for one-component and two-component polariton superfluids in GaAs. The one-component superfluid (dashed black line) has a smaller critical temperature than the two-component superfluid for negative detuning, but larger critical temperature past a certain positive detuning. The two-component superfluid (solid red line) shows a pronounced peak around zero detuning, followed by symmetric decreases for positive and negative detuning values. The parameters used were a Rabi splitting $\Omega_{R}=15$ meV, total polariton concentration $n_{0}=10^{10}$ cm-2, and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
The two-component superfluid exhibits a peak around zero detuning due to balanced contributions from both lower and upper polaritons. Away from resonance, the detuning-dependent Hopfield coefficients redistribute the weights between the polariton branches, leading to symmetric reductions in the critical temperature for the two-component superfluid. These reductions highlight the diminished contributions from interbranch interactions in the two-component superfluid as the system moves further into excitonic or photonic regimes. It is worth noting that for all $\Delta_{0}$ the one-component superfluid has a lower critical temperature than the two-component superfluid. We are also interested in how the critical temperature for our newly developed two-component superfluid compares with the one-component as a function of Rabi splitting in Fig. 9. It is worth commenting about why the two-component superfluid demonstrates a critical temperature for superfluidity that is symmetric around the detuning $\Delta_{0}=0$, while the one-component superfluid demonstrates a peak away from $\Delta_{0}=0$. From Fig. 1, Fig. 6 and Fig. 20 (a), we may arrive at the answer. From Fig. 6 and Eq. (40), we see that the sound velocity is an even function with respect to the detuning. However both the excitonic fraction from Fig. 1 and the effective mass of the lower polariton from Fig. 20 (a) are both not even functions of detuning, and unlike the two-component superfluid, is not cancelled out by symmetric contributions from the upper polariton branch.
Figure 9: Plot of the critical temperature, $T_{c}$ for the one-component and two-component Bose-Einstein condensate, as a function of the Rabi splitting $\Omega_{R}$ in GaAs. The two plots differ in the following way - plot (a) has a negative detuning of $\Delta_{0}=~{}-10$ meV, and plot (b) has a positive detuning of $\Delta_{0}=10$ meV. The parameters used were a total polariton concentration $n_{0}=10^{10}$ cm-2, and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
a) Refer to caption b) Refer to caption
Based on our results in Fig. 9, it is worth restating that for all detuning, the one-component superfluid has a lower critical temperature than the two-component superfluid due to the increased photonic nature of the system, and thus favoring the production of upper polaritons over lower polaritons. We would also like to investigate how the sound velocity and critical temperature vary with the condensate densities, $n_{0}$. In Fig. 10 we plot these results.
Figure 10: Plot of the sound velocity, $c_{s}$ for the one-component and two-polariton condensates, as a function of the total polariton concentration $n_{0}$ in GaAs. The plot represents a positive detuning of $\Delta_{0}=10$ meV. Other parameters used were a Rabi splitting $\Omega_{R}=10$ meV, and a bare photon energy of 1.6 eV.
From Fig. 10, we observe that the sound velocity increases with total polariton concentration. The one-component condensate occupies only the lower polariton branch, with interactions scaled by the square of the excitonic fraction $|X_{P}|^{2}$, and dispersion governed by the effective mass $m_{\text{LP}}$. In contrast, the two-component condensate distributes the same total concentration between both upper and lower polariton branches, with contributions weighted by the Hopfield coefficients. Across the entire concentration range, the two-component condensate exhibits a consistently higher sound velocity, due to the combined effects of both branches and their differing masses and interaction strengths. This emphasizes how including both polariton branches—rather than only one—can enhance collective sound modes in the system. Furthermore, we may investigate how the critical temperature varies with the total polariton concentration. The results are presented in Fig. 11.
Figure 11: Plot of the critical temperature, $T_{c}$ for the one-component and two-polariton superfluids, as a function of the total polariton concentration $n_{0}$ in GaAs. The two plots differ in the following way - plot (a) has a negative detuning of $\Delta_{0}~{}=~{}-10$ meV, and plot (b) has a positive detuning of $\Delta_{0}=10$ meV. The parameters used were a Rabi splitting $\Omega_{R}=10$ meV, and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
a) Refer to caption b) Refer to caption
For both fixed positive and negative detuning, and at fixed Rabi splitting, the one-component superfluid consistently exhibits a lower critical temperature for superfluidity compared to the two-component superfluid. At higher total polariton concentrations, the difference in critical temperature between the one- and two-component polariton superfluids becomes more pronounced, with the two-component superfluid exhibiting a significantly higher $T_{c}$ across the full range.
VI Results for transition metal dichalcogenides
We are also interested in seeing how this analysis holds for transition metal dichalcogenides. In this section, we apply our theoretical framework to TMDCs and plot the results in Fig. 12.
Figure 12: The sound velocity $c_{s}$ as a function of detuning $\Delta_{0}$ for various TMDC materials: MoS2, MoSe2, MoTe2, WS2, and WSe2. The sound velocity is symmetric about $\Delta_{0}=0$, where the detuning corresponds to resonance. TMDCs with lighter exciton effective masses, such as WS2 and WSe2, exhibit higher sound velocities, while heavier effective masses, such as MoTe2, lead to lower sound velocities. The parameters used include a total polariton concentration $n_{0}=10^{12}\,\mathrm{cm}^{-2}$, a bare photon energy of 1.6 eV, and a Rabi splitting $\Omega_{R}=5\,\mathrm{meV}$. The sound velocity is calculated using Eq. (39).
Each curve exhibits a distinctive peak at zero detuning, reflecting the crossover between predominantly excitonic and photonic character in the polariton branches. The magnitude and width of these peaks differ among the materials. Far from zero detuning, the sound velocity approaches a steady value for each TMDC. We may want to examine how the sound velocity, $c_{s}$, changes for chosen detuning, for a range of Rabi splittings. The Rabi splittings can vary widely among TMDCs liu2015 ; dufferwiel2015 ; flatten2016 ; wang2018 and so that would be of immense interest to us. We present the results in Fig.  13.
Figure 13: The sound velocity $c_{s}$ as a function of the Rabi splitting $\Omega_{R}$ for different TMDC monolayers (MoS2, MoSe2, MoTe2, WS2, and WSe2) at a fixed detuning of $\Delta_{0}=10$ meV. The plots demonstrate that the sound velocity increases with increasing Rabi splitting for all TMDCs. The variation between the TMDCs reflects differences in the exciton effective masses, which directly influence the polariton effective masses and interaction strengths. We employ the definition of the sound velocity from Eq. (39).
The results presented in Fig. 13 show the sound velocity $c_{s}$ as a function of the Rabi splitting $\Omega_{R}$ for various TMDC monolayers, specifically MoS2, MoSe2, MoTe2, WS2, and WSe2, for chosen detuning parameter $\Delta_{0}=10.0$ meV. The trend indicates that the sound velocity is increased with increasing Rabi splitting for all these materials. This behavior is due to the fact that the fact that larger Rabi splitting enhances the coupling between the exciton and photon components, leading to changes in the effective polariton masses and interaction strengths. This is in agreement with our results for GaAs in Fig. 5. We may also analyze the critical temperature for superfludity for these TMDCs.
Figure 14: Critical temperature $T_{c}$ as a function of detuning $\Delta_{0}$ for various TMDCs at a fixed Rabi splitting of $\Omega_{R}=8$ meV. The plot illustrates the symmetric dependence of $T_{c}$ on detuning, with maximum occurring at resonance ($\Delta_{0}=0$). The distinct curves for each TMDC highlight the impact of differing exciton effective masses, with lighter effective masses (e.g., WS2 and WSe2) resulting in higher critical temperatures. The parameters used include a total polariton concentration of $n_{0}=10^{10}$ cm-2 and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definition from Eq. (45).
The critical temperature ($T_{c}$) as a function of detuning for different TMDCs is presented in Fig. 15. At zero detuning ($\Delta_{0}=0$), all TMDCs exhibit their maximum $T_{c}$. As the detuning shifts away from zero, the critical temperature decreases symmetrically for both positive and negative detuning values. Notably, TMDCs with lighter exciton effective masses, such as WS2 and WSe2, achieve higher critical temperatures compared to those with heavier exciton masses, such as MoTe2. This highlights the significant influence of material-specific properties in determining the critical temperature, emphasizing the tunability of polaritonic systems for different TMDCs. More importantly, we would like to vary the Rabi splitting for chosen detuning.
Figure 15: Critical temperature $T_{c}$ as a function of Rabi splitting $\Omega_{R}$ for various TMDCs at a fixed detunings of $\Delta_{0}=10$ meV. The plot showcases the increases in critical temperature for fixed detuning of 10 meV. The distinct curves for each TMDC highlight the impact of differing exciton effective masses, with lighter effective masses (e.g., WS2 and WSe2) resulting in higher critical temperatures. The parameters used include a total polariton concentration of $n_{0}=10^{10}$ cm-2 and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
We are also interested in examining how the one-component Bose-Einstein condensate of only lower polaritons, compares with the two-component Bose-Einstein condensate of lower and upper polaritons, when we vary the detuning and the Rabi splitting. The results for the critical temperature versus detuning for this scenario are plotted for four TMDCs of our choice in Fig. 16.
Figure 16: Plots of the critical temperature $T_{c}$ for different TMDCs as a function of the detuning, $\Delta_{0}$ for the two-component condensate and the one-component condensate. Each subfigure corresponds to a specific TMDC - (a) MoS2, (b) MoSe2, (c) WS2, and (d) WSe2. The parameters used are a fixed Rabi splitting $\Omega_{R}=10$ meV, a total polariton concentration of $n_{0}=10^{10}\,\mathrm{cm}^{-2}$ and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
a) [Uncaptioned image] b) [Uncaptioned image]
c) [Uncaptioned image] d) [Uncaptioned image]
We are also interested in how the critical temperature varies with the Rabi splitting in these TMDCs and we plot the results below. From Fig. 16, we observe that the critical temperature as a function of detuning is consistently higher for all selected TMDCs in the two-component superfluid compared to the one-component superfluid consisting solely of lower polaritons, at fixed Rabi splitting and total polariton concentration. At negative detuning, the two-component superfluid shows a significantly higher critical temperature. Although at positive detuning the one-component superfluid eventually surpasses its negative-detuning counterpart in critical temperature, it remains lower than that of the two-component superfluid across the entire range. These results are consistent with our earlier findings for GaAs in Fig. 8.
Figure 17: Plot of the critical temperature, $T_{c}$ for the one-component and two-component Bose-Einstein condensate, as a function of the Rabi splitting $\Omega_{R}$ in WS2. The two plots differ in the following way - plot (a) has a negative detuning of $\Delta_{0}=-10$ meV, and plot (b) has a positive detuning of $\Delta_{0}=10$ meV. The parameters used were a total polariton concentration $n_{0}=10^{10}$ cm-2, and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
a) Refer to caption b) Refer to caption
The results in Fig. 17 are consistent with those obtained for GaAs in Fig. 9. The critical temperature $T_{c}$ for superfluidity is higher in the two-component superfluid of upper and lower polaritons for all fixed negative detunings and varying Rabi splittings $\Omega_{R}$. At fixed positive detuning, the one-component superfluid of lower polaritons yields a higher $T_{c}$ at low Rabi splittings, but as $\Omega_{R}$ increases, the two-component superfluid eventually exhibits a higher critical temperature. In all cases, the total polariton concentration is held fixed. We also investigate how varying the total polariton concentration influences the critical temperature for superfluidity.
Figure 18: Plot of the critical temperature, $T_{c}$ for the one-component and two-polariton condensates, as a function of the total polariton concentration $n_{0}$ in a choice TMDC - MoS2. The two plots differ in the following way - plot (a) has a negative detuning of $\Delta_{0}=-10$ meV, and plot (b) has a positive detuning of $\Delta_{0}=10$ meV. The parameters used were a Rabi splitting $\Omega_{R}=10$ meV, and a bare photon energy of 1.6 eV. The critical temperatures were calculated using the definitions from Eq. (44) and Eq. (45).
a) Refer to caption b) Refer to caption
From Fig. 18, we observe that the two-component superfluid consistently exhibits a higher critical temperature compared to the one-component superfluid across all concentrations and detuning values. This trend mirrors the behavior seen in GaAs, as shown in Fig. 11, where specific choices of detuning and Rabi splitting can lead to a one-component superfluid surpassing the two-component superfluid in critical temperature. While the one-component superfluid may achieve a higher critical temperature under carefully tuned conditions, the two-component superfluid generally maintains a greater critical temperature for superfluidity, particularly at negative detuning and for small positive detuning, in both TMDCs and GaAs.
VII Conclusion and discussion
We have investigated a two-component condensate and superfluid of upper and lower polaritons in semiconductor microcavities and compared our results to those for a one-component condensate and superfluid of lower polaritons only. By retaining the interplay of exciton and photon fractions in both branches, we demonstrated that the effective masses, detuning, and Rabi splitting collectively govern the superfluid behavior of the condensate. Specifically, we have found that the two-component condensate generally exhibits a higher sound velocity and enhanced critical temperature for superfluidity across all detuning. While the one-component superfluid can exhibit a peak in the critical temperature at certain positive detunings, it consistently yields a lower $T_{c}$ compared to the two-component superfluid across the parameter range studied. Increasing the Rabi splitting consistently raises both the sound velocity and critical temperature, underscoring how stronger light–matter coupling can increase the critical temperature for superfluidity.
Our numerical calculations for GaAs quantum wells and various transition metal dichalcogenides (TMDCs) confirm the generality of these trends, indicating the potential of two-component polariton condensates for high-temperature operation across multiple materials. These findings highlight the versatility of multi-branch polariton systems in designing future photonic devices and quantum technologies. By varying the detuning and Rabi splitting, polaritonic devices could utilize the improved stability and tunability offered by two-component condensates. Further theoretical and experimental investigations of spin degrees of freedom, nonlinear interactions, and topological effects may reveal yet richer collective phenomena in exciton–polaritons, paving the way for more advanced applications in coherent light–matter technologies.

### v2

@@ V Superfluidity
where $T_{c}$ is the critical temperature for superfluidity, $\hbar$ is the reduced Planck’s constant, $\rho$ is the total 2D density of the system, accounting for contributions from both polariton branches, $c_{s}$ is the sound velocity defined for the two-component condensate in Eq. (35) and for the one-component condensate are the limits $\alpha=1$ and $\alpha=0$. Additionally, $\zeta(3)$ is the Riemann zeta function evaluated for a value of $3$, and $k_{B}$ is the Boltzmann constant. For the two component condensate of upper and lower polaritons, this is explicitly written in terms of the Hopfield coefficients and effective masses of the upper and lower polartions as follows
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho}{3\zeta(3)k_{B}^{3}}\right]^{1/3}\left(U_{0}{\left(\frac{|X_{0}|^{4}\,n_{\mathrm{LP},0}}{m_{\mathrm{LP}}}\;+\;\frac{|C_{0}|^{4}\,n_{\mathrm{UP},0}}{m_{\mathrm{UP}}}\right)}\right)^{2/3}.$ (40)
We plot the dependance of the critical temperature from Eq. (41) as a function of the detuning, and Rabi splitting.
Figure 5: Critical temperature $T_{c}$ in a two-component polariton condensate in monolayer WSe2. (a) Dependence on detuning $\Delta_{0}$ at fixed Rabi splitting $\Omega_{R}=8$ meV. (b) Dependence on Rabi splitting $\Omega_{R}$ at fixed detuning $\Delta_{0}=5$ meV. In both panels, the total polariton density is held constant at $n_{0}=1\times 10^{11}~\text{cm}^{-2}$, and results are shown for multiple values of the population–split parameter $\alpha$. (c) Contour map of $T_{c}$ as a function of $\Delta_{0}$ and $\Omega_{R}$ at fixed density $n_{0}=1\times 10^{11}~\text{cm}^{-2}$ and population split $\alpha=0.5$ and (d) $\alpha=0.75$. This panel highlights how cavity detuning and light–matter coupling together shape the critical temperature landscape.
At zero detuning ($\Delta_{0}=0$), the lower and upper polariton fractions are equal ($|X_{0}|^{2}=|C_{0}|^{2}=1/2$), so their contributions to Eq. (41) balance, and the critical temperature converges to a single value for all population splits $\alpha$. Away from resonance ($\Delta_{0}\neq 0$), the Hopfield coefficients $|X_{0}|^{2}$ and $|C_{0}|^{2}$ become asymmetric, and the relative condensate populations shift both the sound velocity entering Eq. (41) and the 2D total concentration of the system $\rho$. As shown in Figure 5(a), this causes $T_{c}$ to peak at different detunings depending on $\alpha$. For instance, when the system is predominantly lower polaritonic ($\alpha=0.75$), the enhanced excitonic character increases the interaction-driven sound velocity, which enters as $c_{s}^{4/3}$ in the mean-field expression for $T_{c}$, leading to a higher critical temperature that peaks near $\Delta_{0}\!\approx\!10~\mathrm{meV}$. Increasing the Rabi splitting reduces $T_{c}$ initially for all $\alpha$, before levelling out, with the effect most pronounced for larger values of $\alpha$ where the excitonic contribution to the interaction energy is strongest. In Figure 5, a slightly positive detuning was chosen, so condensates with a greater lower-polariton fraction exhibit higher critical temperatures, reflecting the combined effect of stronger interaction and larger effective mass in determining the superfluid stability.
In addition to the critical temperature, it is convenient to define a critical polariton concentration $n_{c}$ at a given system temperature $T$. This quantity characterizes the minimal density required for superfluidity to emerge at temperature $T$ within our mean-field framework. Starting from the expression for the critical temperature,
$T_{c}=\left[\frac{2\pi\hbar^{2}\rho}{3\zeta(3)k_{B}^{3}}\right]^{1/3}\left[U_{0}\left(\frac{|X_{0}|^{4}\,n_{\mathrm{LP},0}}{m_{\mathrm{LP}}}+\frac{|C_{0}|^{4}\,n_{\mathrm{UP},0}}{m_{\mathrm{UP}}}\right)\right]^{2/3},$ (41)
we substitute the definitions $n_{\mathrm{LP},0}=\alpha n_{0}$ and $n_{\mathrm{UP},0}=(1-\alpha)n_{0}$, and introduce the notation
$S(\Delta_{0},\Omega_{R},\alpha)=\frac{|X_{0}|^{4}}{m_{\mathrm{LP}}}\,\alpha+\frac{|C_{0}|^{4}}{m_{\mathrm{UP}}}\,(1-\alpha).$ (42)
Eq. (41) can then be written as
$T_{c}^{3}=\frac{2\pi\hbar^{2}\rho}{3\zeta(3)k_{B}^{3}}\,\big(U_{0}Sn_{0}\big)^{2}.$ (43)
We also note that the density is defined as,
$\rho=m_{\mathrm{UP}}n_{\mathrm{UP},0}+m_{\mathrm{LP}}n_{\mathrm{LP},0}=n_{0}\Big[m_{\mathrm{LP}}\alpha+m_{\mathrm{UP}}(1-\alpha)\Big]\equiv n_{0}M_{\alpha},$ (44)
where $M_{\alpha}$ is a weighted polariton effective mass determined by the population split parameter $\alpha$. Substituting Eq. (44) for $\rho$ gives
$T_{c}^{3}=\frac{2\pi\hbar^{2}}{3\zeta(3)k_{B}^{3}}\,M_{\alpha}\,(U_{0}S)^{2}\,n_{0}^{3}.$ (45)
Defining
$A\equiv\frac{2\pi\hbar^{2}}{3\zeta(3)k_{B}^{3}},$ (46)
we may write
$T_{c}(\Delta_{0},\Omega_{R},n_{0},\alpha)=\big(AM_{\alpha}\big)^{1/3}\,(U_{0}S)^{2/3}\,n_{0}$} (47)
which shows that $T_{c}$ is linear in the total polariton concentration $n_{0}$ once the density-dependence of $\rho$ is taken into account. Finally, inverting Eq. (47) allows us to identify the critical concentration $n_{c}$ at a given temperature $T$ as
$\boxed{\;n_{c}(T;\Delta_{0},\Omega_{R},\alpha)=\frac{T}{\big(AM_{\alpha}\big)^{1/3}\,(U_{0}S)^{2/3}}\;}.$ (48)
Both $S$ and $M_{\alpha}$ encode the $\alpha$-dependence of the problem. In particular $S$ does so through the Hopfield coefficients and effective polariton masses, and $M_{\alpha}$ through the mass-weighted population imbalance. Thus, measurements of either $T_{c}$ or $n_{c}$ provide a direct probe of the relative occupations of the condensates in the lower and upper polariton branches. The critical concentration $n_{c}$ provides an experimentally accessible quantity that complements the critical temperature $T_{c}$. At fixed detuning $\Delta_{0}$, Rabi splitting $\Omega_{R}$, and interaction strength $U_{0}$, our mean-field framework predicts a one-to-one mapping between $n_{c}$ and the condensate population-split parameter $\alpha$.
This formulation suggests a concrete experimental strategy by measuring the critical density at which superfluidity onsets at a known temperature, an experimentalist can read off the corresponding value of $\alpha$ from our theoretical curves plotted below in Figure (6). This procedure yields two pieces of information simultaneously. The value of $\alpha$, quantifying the relative occupations of the condensates in the lower and upper polariton branches. Additionally it gives confirmation of whether the system indeed realizes a genuine two-component condensate. If the extracted $\alpha$ lies away from the single-component limits ($\alpha=0$ or $\alpha=1$), this serves as direct evidence of coexistence of both branches. Thus, the dependence of $n_{c}$ on $\alpha$ transforms the critical concentration into an experimental probe of population imbalance, offering a pathway to distinguish two-component polariton condensation from its single-component counterparts. An entirely analogous procedure can be applied to the critical temperature $T_{c}$. Measuring $T_{c}$ at fixed density allows one to extract the underlying population imbalance in the same manner.
Figure 6: Condensate population–split parameter $\alpha$ in a two-component polariton condensate in WSe2 monolayers embedded in a GaAs microcavity. (a) Dependence of $\alpha$ on the critical concentration $n_{c}$ at fixed system temperature $T=50$ K, Rabi splitting $\Omega_{R}=8$ meV, and detuning $\Delta_{0}=5$ meV. (b) Dependence of $\alpha$ on the critical temperature $T_{c}$ at fixed total polariton density $n_{0}=1\times 10^{11}~\text{cm}^{-2}$, Rabi splitting $\Omega_{R}=8$ meV, and detuning $\Delta_{0}=10$ meV.
VI Discussion and conclusions
We have developed a microscopic mean-field theory of two-component superfluidity in semiconductor microcavities, treating coexisting condensates of lower and upper polaritons as coupled components of a unified superfluid, with their distinct dispersions, effective masses, and interaction strengths emerging from the underlying exciton-photon Hamiltonian. By incorporating the full momentum and detuning dependence of the Hopfield coefficients, effective masses, and interaction strengths, we derived explicit expressions for the collective excitation spectrum, sound velocity $c_{s}$, and critical temperature $T_{c}$ as functions of the cavity detuning $\Delta_{0}$, Rabi splitting $\Omega_{R}$, and our newly introduced, phenomenological population-split parameter $\alpha$. Our central result is that two-component condensation produces experimentally measurable signatures that distinguish it from conventional single-branch superfluidity, that is enhanced sound velocity and critical temperature across most of the parameter space, distinctive detuning-dependent peak structures in $c_{s}(\Delta_{0},\Omega_{R})$, $T_{c}(\Delta_{0},\Omega_{R})$, and $n_{c}(\Delta_{0},\Omega_{R})$ and a systematic dependence on $\alpha$ that enables direct experimental extraction of the population split from measurements at known system parameters.
At zero detuning, where $|X_{0}|^{2}=|C_{0}|^{2}=1/2$, the system is symmetric and $c_{s}$ and $T_{c}$ converge to single values independent of $\alpha$. Away from resonance, the Hopfield-weighted interaction terms $g_{\ell\ell}=U_{0}|X_{0}|^{4}$, $g_{uu}=U_{0}|C_{0}|^{4}$, and $g_{\ell u}=U_{0}|X_{0}|^{2}|C_{0}|^{2}$ become asymmetric, and both $c_{s}$ and $T_{c}$ develop pronounced $\alpha$-dependence. For instance, at positive detunings in monolayer WSe2 ($\Delta_{0}=5~\mathrm{meV}$, $\Omega_{R}=8~\mathrm{meV}$), an LP-dominated condensate ($\alpha=0.75$) reaches a peak critical temperature $T_{c}\approx 140~\mathrm{K}$, compared to $T_{c}\approx 120~\mathrm{K}$ for a balanced condensate ($\alpha=0.5$). These enhancements arise from the stronger excitonic contribution to the interaction energy when the LP branch is predominantly excitonic in nature.
Many examples demonstrate that multi-component condensation with interbranch coherence is not merely hypothetical but has been realized experimentally in systems with comparable energy splittings and interaction-driven coupling. The magnon case is particularly relevant: Bailey et al. (2024) Aeppli2022 report multi-band Bose-Einstein condensation stabilized by four-particle scattering processes, demonstrating that interaction-mediated coupling can sustain coexisting populations in energetically distinct branches. This experiment was performed in thin films of Yttrium Iron Garnet (YIG), where quantized spin waves (magnons) form quantized subbands due to confinement across the film thickness. Under microwave pumping, Bose–Einstein condensation occurring simultaneously in two distinct magnon bands has been observed. One corresponds to the lowest (Kittel-like) mode and the other to an excited quantized mode. The coexistence of these two condensates is stabilized by strong four-magnon scattering processes that exchange pairs of magnons between the bands, effectively coupling the condensates despite their energy splitting. This establishes a genuine multi-band condensate with interaction-mediated coherence between energetically distinct branches. It is directly analogous to our LP–UP system, where interbranch scattering terms in Eq. (15) play the same coupling role, suggesting that two-component condensation stabilized by interaction-driven coupling is experimentally realistic.
Crucially, our framework is intentionally predictive rather than explanatory. We do not derive the microscopic conditions under which simultaneous LP and UP condensation occurs—that question requires a full kinetic theory incorporating reservoir dynamics, phonon scattering, pump mechanisms, and the conditions for phase-locking between branches—and remains an important target for future work. Instead, we establish theoretical benchmarks that answer the question: if both branches are macroscopically occupied, what collective phenomena emerge, and how can they be unambiguously identified?
Figure 6 provide explicit extraction protocols in that by measuring $c_{s}$ or $T_{c}$ at known $n_{0}$ and $\Delta_{0}$ and $\Omega_{R}$, experimentalists can infer $\alpha$ and thereby confirm the presence and degree of two-component condensation. These predictions are general and apply to any strongly coupled exciton-photon system, with numerical parameters adjusted for specific materials (GaAs/AlGaAs quantum wells and monolayer TMDCs embedded in GaAs microcavities).
While most experiments to date observe single-branch condensation due to efficient phonon-assisted relaxation, recent demonstrations of upper-polariton condensation Chen2023 , suggest that controlled dual-branch occupation may be experimentally accessible under appropriate conditions. These developments echo successful realizations of multi-component condensation in related systems: spinor atomic BECs exhibit simultaneous occupation of multiple hyperfine states with interaction-driven phase-locking Stamper2013 ; Ho1998 , magnon condensates display coexisting branches with hybridized collective modes Aeppli2022 . Our results provide the theoretical framework needed to recognize and characterize analogous phenomena when LP and UP branches are simultaneously macroscopically occupied.
Looking forward, several natural extensions of this work merit investigation. A complete understanding of the coupled LP–UP condensate will ultimately require a time-dependent treatment of the relative phase between the two order parameters, which may give rise to Rabi-like oscillations or other dynamical coherence phenomena beyond the scope of the present equilibrium model. First, a driven–dissipative formulation based on the Lindblad master equation would capture the intrinsically nonequilibrium nature of polariton condensates, incorporating pumping, decay, and reservoir dynamics explicitly to provide time-dependent predictions for $\alpha(t)$ and to determine the conditions under which steady-state dual-branch occupation can emerge. Second, a full amplitude–phase decomposition of the coupled order parameters—together with a microscopic analysis of the phase-locking condition—remains an essential next step. Such a study would establish how interaction-induced coherence and fluctuations determine the stability of the relative phase, thereby connecting the phenomenological population-split parameter $\alpha$ to experimentally controllable pump and reservoir parameters. Finally, our framework naturally generalizes to other multi-component light–matter systems, including cavity magnon-polaritons, phonon-polaritons, and multi-mode photonic condensates, where interaction-driven coupling between branches may yield analogous collective phenomena. By formulating a quantitative equilibrium baseline for two-component polariton superfluidity, this work provides both conceptual motivation and measurable benchmarks for future microscopic and experimental investigations.